OpenStudy (anonymous):

\[\sqrt{3}+\sqrt{75}\]

OpenStudy (anonymous):

\[\sqrt{3}+\sqrt{75}= \sqrt{3}+ 5\sqrt{3} = 6\sqrt{?}\]

OpenStudy (anonymous):

3

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

\[6\sqrt{3}\]

OpenStudy (anonymous):

If I deserve medal then give me one

OpenStudy (anonymous):

Yes. I don't know what's up with Latex today...

OpenStudy (anonymous):

T Y

myininaya (myininaya):

lol so that wasn't suppose to be fill in the blank lol?

OpenStudy (anonymous):

guess not

OpenStudy (anonymous):

Yeah... just a bad Latex error :D